OpenStudy (anonymous):

Which is the correct simplification of 5.4x^12/1.2x10^3 ? confused! help please

OpenStudy (amistre64):

i think your missing a 10 in that

OpenStudy (amistre64):

5.4 x 10^12 ---------- 1.2 x 10^3 by chance?

OpenStudy (anonymous):

yes your correct, i meant 5.4x10^12

OpenStudy (amistre64):

well, its good to know that: \[\frac{b^n}{b^m}=b^{n-m}\]

OpenStudy (anonymous):

Wow i got a Cac Error lol There we go :p Crap messed up and i think not lol

OpenStudy (amistre64):

5.4 x 10^12 5.4 x 10^(12-3) ---------- = -------------- 1.2 x 10^3 1.2

OpenStudy (anonymous):

Thats what i got

OpenStudy (amistre64):

54 x 10^(9) --------- 12 27 x 10^(9) --------- 6 9 x 10^(9) --------- ; so far :) 2

OpenStudy (amistre64):

9/2 = 4.5 soooo 4.5 x 10^9 is my best guess

OpenStudy (anonymous):

thank you! i dont think i ever would have gottn it :P

OpenStudy (amistre64):

yw